# HybridMQA: Exploring Geometry-Texture Interactions for Colored Mesh Quality Assessment

### v1

@@ Abstract
Mesh quality assessment (MQA) models play a critical role in the design, optimization, and evaluation of mesh operation systems in a wide variety of applications. Current MQA models, whether model-based methods using topology-aware features or projection-based approaches working on rendered 2D projections, often fail to capture the intricate interactions between texture and 3D geometry. We introduce HybridMQA, a first-of-its-kind hybrid full-reference colored MQA framework that integrates model-based and projection-based approaches, capturing complex interactions between textural information and 3D structures for enriched quality representations. Our method employs graph learning to extract detailed 3D representations, which are then projected to 2D using a novel feature rendering process that precisely aligns them with colored projections. This enables the exploration of geometry-texture interactions via cross-attention, producing comprehensive mesh quality representations. Extensive experiments demonstrate HybridMQA’s superior performance across diverse datasets, highlighting its ability to effectively leverage geometry-texture interactions for a thorough understanding of mesh quality. Our implementation will be made publicly available.
@@ 1 Introduction
Advancements in 3D capture and display technologies have sparked a growing interest in immersive media. 3D meshes, a key form of 3D media, are widely used in applications like virtual and augmented reality [21, 51], gaming, animation, medical modeling [39], and generative 3D content creation [22, 37]. A mesh, comprising triangular faces formed by vertices, is colorized by assigning RGB colors to each vertex (vertex-color mesh) or applying a 2D texture map with UV coordinates (textured mesh). The demand for colored meshes calls for high-quality acquisition [3], compression [11, 26], and transmission [27]. However, these processes often introduce artifacts that degrade visual quality, highlighting the need for robust mesh quality assessment (MQA) methods. Full-reference (FR) approaches, which compare distorted meshes against pristine references, are essential for accurate visual quality assessment.
@@ 3.2 Graph Convolutional Network
Here, $\bm{f}_{i}\in\mathbb{R}^{C_{1}}$ is the sampled feature for the $i$-th vertex, $u^{h}_{i}$ and $u^{w}_{i}$ are its UV coordinates, and $\bm{T}_{BE}$ denotes base encoder’s output feature maps. We average the sampled features for vertices with multiple UV mappings. This process produces a feature graph with vertices initialized by sampled features and edges defined by mesh connectivity $\bm{E}$.
@@ 3.4 Quality Encoder
We design a quality encoder that processes the aligned projections from model and texture branches to output an expressive quality representation for the mesh. It captures the geometry-texture interactions between the two sets of projections through multi-scale cross-attentions. As discussed in Sec. 1, understanding these interactions is key to a robust and effective MQA. Given model branch’s 3D feature projections $\bm{I}_{m}$ and texture branch’s colored projections $\bm{I}_{t}$, the quality encoder $\psi$ outputs
@@
where $\bm{f}_{mesh}$ denotes the final quality representation, and $\bm{\theta}_{\psi}$ denotes the learnable parameters. Figure 4(a) illustrates the process in detail.
@@
Cross-attention Modules. We apply cross-attention to capture the interactions between the aligned 3D surface representations $\bm{f}^{s}_{m,p}$ and color representations $\bm{f}^{s}_{t,p}$ at each scale. Figure 4(b) depicts our proposed module with two transformer blocks [42], where we alternate the query and key-value roles of $\bm{f}^{s}_{m,p}$ and $\bm{f}^{s}_{t,p}$ to explore their interactions and mutual influence. This simulates how one domain’s representation affects the impact of the other domain’s representation on perceptual quality, as discussed in Sec. 1. We concatenate the outputs to form the single-scale quality representation $\bm{f}^{s}_{p}$.
@@ 4.1 Datasets & Implementation Details
We use ResNet50 [14] pre-trained on ImageNet [8] as the image encoder $\varphi$ and a CNN with randomly initialized weights as the 3D feature encoder $\phi$. 3D feature and colored projections are rendered at $128\times 128$ and $512\times 512$ resolutions, with patch sizes of $16\times 16$ and $64\times 64$, respectively, to ensure alignment. For data augmentation, we randomly perturb camera angles and flip the patches in training to improve robustness and generalization. More details of the datasets and implementation are provided in the supplementary material.
@@ 4.3 Quantitative Results
LPIPS 0.592 0.584 0.711 0.695
Graphics-LPIPS 0.731 0.734 0.801 0.795
HybridMQA 0.800 0.783 0.857 0.825
@@ 4.5 Ablation Studies
exc. Texture map 0.866 0.872
exc. Normal map 0.863 0.866
exc. Vertex map 0.872 0.873
exc. Base Encoder 0.856 0.851
exc. GCN 0.866 0.865
all (proposed) 0.892 0.897
@@ 5 Conclusion & Discussion
A few limitations present room for improvement. Our method relies on perpendicular viewpoints without considering their varying contributions to perceptual quality. Integrating a semantic-aware module could address this by weighting viewpoints based on importance. Also, memory consumption scales with mesh size as the GCN processes the entire mesh graph, limiting efficiency in real-world applications. Sampling techniques [52, 17, 53] could help reduce memory footprint. Future directions include adapting our work for no-reference MQA, point cloud quality assessment, and exploring applications in perceptually optimized mesh compression, enhancement, and generation.

### v2

@@ Abstract
Mesh quality assessment (MQA) models play a critical role in the design, optimization, and evaluation of mesh operation systems in a wide variety of applications. Current MQA models, whether model-based methods using topology-aware features or projection-based approaches working on rendered 2D projections, often fail to capture the intricate interactions between texture and 3D geometry. We introduce HybridMQA, a first-of-its-kind hybrid full-reference colored MQA framework that integrates model-based and projection-based approaches, capturing complex interactions between textural information and 3D structures for enriched quality representations. Our method employs graph learning to extract detailed 3D representations, which are then projected to 2D using a novel feature rendering process that precisely aligns them with colored projections. This enables the exploration of geometry-texture interactions via cross-attention, producing comprehensive mesh quality representations. Extensive experiments demonstrate HybridMQA’s superior performance across diverse datasets, highlighting its ability to effectively leverage geometry-texture interactions for a thorough understanding of mesh quality. Our project website is available at https://arshafiee.github.io/hybridmqa/.
@@ 1 Introduction
Advancements in 3D capture and display technologies have sparked a growing interest in immersive media. 3D meshes, a key form of 3D media, are widely used in applications like virtual and augmented reality [21, 51], gaming, animation, medical modeling [39], and generative 3D content creation [22, 37]. A mesh, comprising triangular faces formed by vertices, is colorized by assigning RGB colors to each vertex (vertex-color mesh) or applying a 2D texture map with UV coordinates (textured mesh). The demand for colored meshes calls for high-quality acquisition [3], compression [11, 26], and transmission [27]. However, these processes often introduce artifacts that degrade visual quality, highlighting the need for robust mesh quality assessment (MQA) methods. Full-reference (FR) approaches, which take distorted meshes and their pristine references as input and generate a quality score by comparing their visual quality, are essential for accurate mesh quality assessment.
@@ 3.2 Graph Convolutional Network
Here, $\bm{f}_{i}\in\mathbb{R}^{C_{1}}$ is the sampled feature for $i$-th vertex, $u^{h}_{i}$ and $u^{w}_{i}$ are its UV coordinates, and $\bm{T}_{BE}$ denotes base encoder’s output feature maps. For vertices with multiple UV mappings, we average the sampled features. This process produces a feature graph with vertices initialized by sampled features and edges defined by mesh connectivity $\bm{E}$.
@@ 3.4 Quality Encoder
We design a quality encoder that processes the aligned projections from model and texture branches to output an expressive quality representation for the mesh. It captures the geometry-texture interactions between the two sets of projections through multi-scale cross-attentions. As discussed in Sec. 1, understanding these interactions is key to a robust and effective MQA. As detailed in Fig. 4(a), given model branch’s 3D feature projections $\bm{I}_{m}$ and texture branch’s colored projections $\bm{I}_{t}$, the quality encoder $\psi$ outputs
@@
where $\bm{f}_{mesh}$ denotes the final quality representation, and $\bm{\theta}_{\psi}$ denotes the learnable parameters.
@@
Cross-attention Modules. We apply cross-attention to capture the interactions between aligned 3D surface representations $\bm{f}^{s}_{m,p}$ and color representations $\bm{f}^{s}_{t,p}$ at each scale. Figure 4(b) depicts our proposed module with two transformer blocks [42], where we alternate the query and key-value roles of $\bm{f}^{s}_{m,p}$ and $\bm{f}^{s}_{t,p}$ to explore their interactions and mutual influence. This simulates how one domain’s representation affects the impact of the other domain’s representation on perceptual quality, as discussed in Sec. 1. We concatenate outputs to form single-scale quality representation $\bm{f}^{s}_{p}$.
@@ 4.1 Datasets & Implementation Details
We use ResNet50 [14] pre-trained on ImageNet [8] as the image encoder $\varphi$ and a randomly initialized CNN as the 3D feature encoder $\phi$. 3D feature and colored projections are rendered at $128\times 128$ and $512\times 512$ resolutions, with patch sizes of $16\times 16$ and $64\times 64$, respectively, to ensure alignment. For data augmentation, we randomly perturb camera angles and flip the patches in training to improve robustness and generalization. More details of the datasets and implementation are provided in the supplementary material.
@@ 4.3 Quantitative Results
LPIPS 0.592 0.584 0.712 0.695
Graphics-LPIPS 0.731 0.734 0.784 0.773
HybridMQA 0.800 0.783 0.846 0.811
@@ 4.5 Ablation Studies
w/o Texture map 0.866 0.872
w/o Normal map 0.863 0.866
w/o Vertex map 0.872 0.873
w/o Base Encoder 0.856 0.851
w/o GCN 0.866 0.865
HybridMQA (proposed) 0.892 0.897
@@ 5 Conclusion & Discussion
A few limitations present room for improvement. Our method relies on perpendicular viewpoints without considering their varying contributions to perceptual quality. Integrating a semantic-aware module could address this by weighting viewpoints based on importance. Also, memory consumption scales with mesh size as the GCN processes the entire mesh graph, limiting efficiency in real-world applications. Sampling techniques [52, 17, 53] could help reduce memory footprint. Future directions include (1) adapting our work for no-reference MQA; (2) generalizing it to point cloud quality assessment [56, 58] by defining graph edges; and (3) exploring applications in perceptually optimized mesh compression, enhancement, and generation.
[…]
Supplementary Material
6 Experimental Setup Details
6.1 Details of Datasets
To validate the performance of our proposed method, we conduct experiments on four publicly available color MQA datasets: Nehmé et al. [29], SJTU-TMQA [7], TSMD [49], and CMDM [30]. The Nehmé et al. dataset is the largest public dataset of 3D textured meshes, containing 55 source meshes distorted by a mixture of geometric and color distortions to obtain 3000 distorted meshes. The SJTU-TMQA dataset consists of 21 reference and 945 distorted textured meshes. Distorted meshes were generated through geometric or color distortions or a combination of both. The TSMD dataset includes 39 source 3D textured meshes (excluding 3 source meshes as they were not publicly available: “Mitch", “Nathalie", and “Thomas"), each distorted at five levels with a combination of geometric and color distortions, resulting in a total of 195 distorted meshes. Finally, the CMDM dataset consists of vertex-color meshes, with 5 source meshes each subjected to geometric or color distortions, resulting in 80 distorted meshes. Mean opinion scores (MOS) were computed and reported as ground truth quality labels for all distorted models across the four datasets, based on subjective evaluations from 4513, 73, 74, and 72 study participants, respectively. In total, the four datasets encompass a wide variety and strength levels of geometric and color distortions. We note that the TSMD and SJTU-TMQA datasets have overlapping source meshes which were excluded from the training set (TSMD dataset) in our generalization test.
6.2 Implementation Details
We use Adam optimizer [15] with the default $1e^{-5}$ weight decay and $1e^{-4}$ initial learning rate that is gradually reduced to $1e^{-5}$ with cosine annealing scheduler [24]. The default batch size is set to 8, and the model is trained for 15 epochs by default. The loss balance term $\lambda$ is set to 1. During training and testing on the CMDM dataset, we skip the base encoder and directly initialize the feature graph with raw vertex color, normal, and position values as vertex-color meshes lack 2D texture maps and UV mapping data. To allow for faster training and larger batch sizes given the limitations of our GPU (NVIDIA V100 32GB), we implement viewpoint dropout, where we randomly select two out of six camera viewpoints in each training iteration and only render those two projections.
Data Augmentation. We use camera angle augmentation in training to enhance the model’s robustness and generalization capabilities. Specifically, we set the original azimuth and elevation angles as the mean of a normal distribution with a standard deviation of $22.5^{\circ}$ and sample new azimuth and elevation angles in each training iteration. We also employ flip augmentation on patches extracted from 3D feature and colored projections.
6.3 Details of Evaluation Metrics
To compare the performance of different MQA methods, we employ two mainstream evaluation criteria: the Spearman rank-order correlation coefficient (SRCC) and the Pearson linear correlation coefficient (PLCC). SRCC measures prediction monotonicity, while PLCC evaluates prediction accuracy [12]. The PLCC score is calculated by using a logistic non-linear fitting method to align the predicted scores with the ground truth scale [12]. Higher SRCC and PLCC absolute values signal a higher correlation between MOS and predicted quality scores and hence a better performance.
7 Further Ablation Studies
We perform additional ablation experiments on Nehmé et al. dataset [29].
7.1 Cross-attention Mechanism
We perform further ablation studies to highlight the impact of the cross-attention mechanism. Specifically, given the encoded 3D surface representation $\bm{f}_{m}$ and the textural representation $\bm{f}_{t}$, we replace the proposed cross-attention mechanism with: (1) addition; (2) weighted addition of $\bm{f}_{m}$ and $\bm{f}_{t}$, where we learn the weights using a convolutional block that takes the two representations as input; (3) concatenation; (4) elementwise multiplication; and (5) self-attention of $\bm{f}_{m}$ and $\bm{f}_{t}$ followed by concatenation. Table 5 presents the results. We can observe that all replacements result in significant drops in performance. This highlights the effectiveness of the proposed cross-attention mechanism in capturing interactions between 3D geometry and textural representations of the mesh, emphasizing the importance of these texture-geometry interactions for achieving accurate MQA.
Figure 8: HybridMQA clearly outperforms Graphics-LPIPS [29] in gMAD competition [25]. Columns one and two showcase results with Graphics-LPIPS fixed at low and high quality, respectively, while columns three and four display results with HybridMQA fixed at low and high quality. In each column, the left objects are the references, while the right ones are the distorted meshes. The most perceptually important viewpoint of each object is selected for visualization.
Configurations SRCC PLCC
addition: $\bm{f}_{m}+\bm{f}_{t}$ 0.842 0.842
weighted addition: $\bm{f}_{m}+\bm{w}\odot\bm{f}_{t}$ 0.846 0.861
concat.: $\bm{f}_{m}\oplus\bm{f}_{t}$ 0.845 0.857
multiplication: $\bm{f}_{m}\odot\bm{f}_{t}$ 0.848 0.849
self-att. + concat.: $\text{SA}(\bm{f}_{m})\oplus\text{SA}(\bm{f}_{t})$ 0.852 0.857
cross-attention (proposed) 0.892 0.897
Table 5: Ablation on cross-attention mechanism on Nehmé et al.
7.2 Data Augmentations
We also conduct experiments to measure the importance of camera angle and flip augmentations in the method’s performance. Table 6 presents the results of excluding each of the two data augmentations. We observe that both data augmentations improve performance, with camera angle augmentation having a more pronounced effect.
Angle Aug. Flip Aug. SRCC PLCC
0.876 0.883
0.857 0.857
0.892 0.897
Table 6: Ablation on data augmentations on Nehmé et al.
7.3 Viewpoint Dropout & Batch Size
We conduct further experiments to evaluate different configurations of viewpoint dropout and batch size, as introduced in Sec. 6.2. Specifically, we evaluate three configurations: randomly selecting two or four viewpoints in each training iteration or using all six viewpoints (no dropout). These configurations are tested across batch sizes of 2, 4, and 8. We note that the largest possible batch size varies depending on the number of viewpoints: 8 for two viewpoints, 4 for four viewpoints, and 2 for six viewpoints. Table 7 presents the results. We can see that performance improves as the batch size increases for each viewpoint configuration. Notably, the best performance is achieved with two viewpoints, which allows for a batch size of 8—the largest among the tested configurations. This demonstrates the effectiveness of the viewpoint dropout mechanism.
$N_{v}\backslash N_{b}$ 2 4 8
2 Views 0.837/0.844 0.864/0.873 0.892/0.897
4 Views 0.859/0.867 0.866/0.873 OOM
6 Views 0.838/0.846 OOM OOM
Table 7: SRCC/PLCC results of the ablation on the number of viewpoints and batch sizes in training on Nehmé et al. $N_{v}$ and $N_{b}$ denote the number of viewpoints and batch size, respectively. OOM stands for out of memory.
8 Further Qualitative Results
8.1 gMAD Competition
We perform gMAD competition [25] to qualitatively compare the performance of HybridMQA with Graphics-LPIPS [29]. gMAD competition identifies 3D meshes that one method estimates to be of similar quality, while the other method rates them as having significantly different quality. Through this competition, at least one of the methods will be discredited due to producing quality judgments that do not correlate with human opinions. We perform the gMAD competition on the SJTU-TMQA dataset [7], where we gather quality judgments of the two methods on all validation sets of the 5-fold cross-validation test.
Figure 8 presents the results of the competition, where HybridMQA clearly outperforms Graphics-LPIPS. As we can see, Graphics-LPIPS judges the 3D meshes in the first column (pottery vessel and watermelon) to be of similarly low quality. This is clearly in contradiction with human judgments as well as HybridMQA predictions. The second column shows a similar trend: HybridMQA predictions align with human judgments, while Graphics-LPIPS incorrectly rates the girl 3D mesh as having high quality. We then switch the roles of the two methods in the third and fourth columns. In column three, Graphics-LPIPS assigns higher quality prediction to the girl compared to the bread. However, both 3D meshes are severely contaminated by JPEG compression [7] and judged by human viewers to be of similarly low quality. HybridMQA successfully rates the two meshes as having poor perceptual quality. Similar conclusions can be made in the fourth column, where HybridMQA accurately assigns high quality scores to both meshes. These results demonstrate the clear superiority of HybridMQA over Graphics-LPIPS in colored MQA.
8.2 GradCAM on meshes
Figure 9: More GradCAM [35] results on meshes.
Figure 10: More GradCAM [35] results on meshes.
Figures 9 and 10 provide additional examples of GradCAM [35] applied to graph features in the model branch. The highlighted regions successfully identify noticeable geometrical artifacts that align well with human perception. This showcases the model branch’s effectiveness in capturing geometry-aware quality representations.
8.3 GradCAM on Cross-attention
Figure 11 provides additional examples of GradCAM [35] applied before and after cross-attention. The two branches concentrate on distinct regions, with the model branch emphasizing geometric artifacts. Through cross-attention, the framework effectively identifies and focuses on perceptually important regions by exploring interactions between geometry and texture. This demonstrates the effectiveness of our hybrid method in exploiting interactions between representations learned in texture and model branches.
Figure 11: More GradCAM [35] results on cross-attention.